OpenStudy (anonymous):

Please Simplify. Show your work. 5 1/3 +(-3 9/18)

OpenStudy (cj49):

5 1/3= 5*3+1/3=16/3 -3 9/18=-3*18+9/18=63/18=21/6=7/2 16/3-7/2 (16*2)-(7*3)/6=?

jhonyy9 (jhonyy9):

so first of all 9/18 = 1/2

jhonyy9 (jhonyy9):

and then -3 1/2 = -7/2

jhonyy9 (jhonyy9):

so then 16/3 + (-7/2) = 16/3 -7/2 = ((16*2) -(7*3))/6 = (32-21)/6 = 11/6

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

I believe so.

jhonyy9 (jhonyy9):

ok was my pleasure bye